OpenStudy (anonymous):

Given a dilation with the origin O (0, 0), by observation determine the scale factor "K." D o,k = (6, 12) ---> (3, 6)

jimthompson5910 (jim_thompson5910):

hint: multiply the coordinates of (6,12) by k to get (3,6)

OpenStudy (anonymous):

is my answer-3?

jimthompson5910 (jim_thompson5910):

how are you getting -3

OpenStudy (anonymous):

wait. thats not right. i dont know how I subtracted that. Meant to say is my answer 2

OpenStudy (anonymous):

only other answer I came up with

jimthompson5910 (jim_thompson5910):

how are you getting 2 now?

OpenStudy (anonymous):

2 goes into 3, 6 an can make 12.

jimthompson5910 (jim_thompson5910):

so to go from (6,12) to (3,6) are you a) multiplying both coordinates by 2 OR b) dividing both coordinates by 2 ?

OpenStudy (anonymous):

divide

jimthompson5910 (jim_thompson5910):

but you said that k = 2

jimthompson5910 (jim_thompson5910):

if k = 2, that means that you're multiplying the coordinates by 2

OpenStudy (anonymous):

Honestly I dont know what I am doing. I am confusing myself lol

jimthompson5910 (jim_thompson5910):

the first point is (6,12) you multiply each coordinate by some number k to get (3,6)

jimthompson5910 (jim_thompson5910):

so... 6*k = 3 and 12*k = 6

OpenStudy (anonymous):

Oh. . that makes sense now lol

OpenStudy (anonymous):

Thats why theres a 3/6. but what about finding the final answer?

jimthompson5910 (jim_thompson5910):

3/6 isn't fully reduced though

OpenStudy (anonymous):

Hmm

jimthompson5910 (jim_thompson5910):

6*k = 3 6*k/6 = 3/6 ... Divide both sides by 6 to isolate k. k = ???

OpenStudy (anonymous):

1/2?

jimthompson5910 (jim_thompson5910):

good, k = 1/2